OpenStudy (anonymous):

the quantity of x plus 20 over 2 = 3x

OpenStudy (anonymous):

x + 20/2 = 3x multiple both sides by 2 to get rid of the 2 2(x+20/2) = 3x(2) the 2s cancel each other out on the left side x+20=6x you want to get the x's on the same side so subtract x from each side 20=5x divide both sides by 5 to get x on its own 20/5 = 5x/5 4=x

OpenStudy (anonymous):

x + 20/2 = 3x or 2x=10 x=5

OpenStudy (anonymous):

ty whats the answr to this? 1 over 3 (2x − 8) = 4

OpenStudy (anonymous):

@doubledrive what aout this 1 over 3 (2x − 8) = 4

OpenStudy (anonymous):

Multiply both sides of the equation by 3 to cancel the fraction 3((1/3)(2x-8)) = 4(3) 1(2x-8) = 12 2x - 8 = 12 2x = 12 + 8 2x = 20 x = 20/2 x = 10

OpenStudy (anonymous):

ty

OpenStudy (valpey):

You sure you aren't asking?: \[\frac{(x+20)}{2}= 3x\]